A cubic solid is made of two elements P and Q. All P cations are located at the corners of the unit cell and six Q anions on the center of all the cubic faces.
i. Draw a structure of the cubic mentioned above.
ii. What is the coordination number of Q?
iii. What is the formula of the compound? Show your work.
1
Expert's answer
2020-09-24T08:13:28-0400

( i )

Structure in which  P cations are located at the corners of the unit cell and six Q anions on the center of all the cubic faces.



( ii )

 coordination number of Q is " 12 " due to fcc arrangement .


( iii )

 formula of the compound is PQ3 .



Explanation =


effective number of atoms of p = 1/8 * 8 = 1 ( due to all corners )



effective number of atoms of q = 1/2 * 6 = 3 ( due to all face center )